# REACT: Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation

### v2

@@ Abstract
Scene Graph Generation (SGG) is a task that encodes visual relationships between objects in images as graph structures. SGG shows significant promise as a foundational component for downstream tasks, such as reasoning for embodied agents. To enable real-time applications, SGG must address the trade-off between performance and inference speed. However, current methods tend to focus on one of the following: (1) improving relation prediction accuracy, (2) enhancing object detection accuracy, or (3) reducing latency, without aiming to balance all three objectives simultaneously. To address this limitation, we propose a novel architecture, inference method, and relation prediction model. Our proposed solution, the REACT model, achieves the highest inference speed among existing SGG models, improving object detection accuracy without sacrificing relation prediction performance. Compared to state-of-the-art approaches, REACT is 2.7 times faster (with a latency of 23 ms) and improves object detection accuracy by 58.51%. Furthermore, our proposal significantly reduces model size, with an average of 5.5x fewer parameters [^1] Code is available at https://github.com/Maelic/SGG-Benchmark.
@@ 1 Introduction
Figure 1: Performance comparison between our approach and previous work, in terms of performance (F1@K metric) and Latency (ms) on the PSG dataset [42]. The points are scaled by the number of parameters for each model (smaller is better).
Scene Graph Generation (SGG) is the task of generating a structured graph representation from visual inputs. Given an image, visual regions corresponding to objects of interest are extracted, and relations between these objects are predicted in the form of $<subject,predicate,object>$ triplets. Aggregated together, these triplets form a directed acyclic graph which can be interpreted as a structured representation of the scene. In addition, subject and object nodes are traditionally associated with their coordinates on the image (i.e., bounding boxes). In recent years, this task has garnered increasing attention thanks to its utility in various downstream applications, such as Visual Question Answering (VQA) [8, 5, 23], Image Captioning [43, 40, 30], or Reasoning of an embodied agent [7, 22, 28]. However, we observe a significant unbalance between the progress in the SGG task itself and its adoption as a foundational component for downstream tasks [2]. One of the reasons is the lack of low-cost approaches that could power applications with real-time constraints. Because SGG is a task that combines Object Detection (OD) and Relation Prediction (RelPred), a good trade-off between accuracy in these two parameters is also required. Nonetheless, current approaches focus on one or two of these parameters without considering the general problem. In this work, we explore the bottlenecks in SGG to find an optimal combination of OD accuracy, RelPred accuracy, and latency (see Figure 1).
We can separate SGG methods into two categories: Two-Stage approaches (TS) and One-Stage approaches (OS). The former uses a combination of an object detector and a relation predictor to model objects and relations in a sequence. The latter uses a single-stage pipeline, to infer relations and object proposals directly from the image features. As a result, one-stage approaches are often more efficient for real-time processing [10]. However, a few challenges remain, especially concerning the accuracy of such approaches in the OD task [20, 9]. In contrast, two-stage approaches train the object detector and relation predictor sequentially, leading to improved performance in object detection.
In this work, we propose to investigate the bottlenecks of the two-stage architecture, in terms of OD accuracy and latency. First, structural dependencies between blocks of the architecture are reviewed as a Decoupled Two-Stage SGG (DTS-SGG). This design permits to maintain performance of the pre-trained detector and allows better control of the information flow between the two stages of the architecture. DTS enables the selection of a smaller set of high-quality objects (i.e., proposals) before passing them to the relation prediction stage. This approach reduces overall latency, as the complexity of relation prediction increases exponentially with the number of input objects. To select an optimal amount of proposals at inference time, we define a new algorithm, the Dynamic Candidate Selection (DCS). DCS uses evaluation results to find the best trade-off between latency and accuracy at test time. Thanks to the DTS architecture, we can switch the dependency on a Faster-RCNN model for any real-time object detectors, such as one of the YOLO family (YOLOV8/9/10 [15, 39, 38]). Lastly, popular architectures in SGG [46] use a combination of different feature sources, such as union pair features [44] or textual features of the subject and object classes [27]. This results in a significant amount of processing to gather and refine each type of feature altogether. By removing the dependence on sub-optimal features, we can reduce the parameter count of the model and lower the latency. By coupling this strategy to the DTS architecture, we propose a new model for SGG, the Real-Time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation (REACT). The contributions of this work are three-fold:
  1. 1.
    A new architecture for SGG, the Decoupled Two-Stage (DTS). Thanks to this new architecture, real-time object detectors can be coupled with state-of-the-art two-stage RelPred components, and lead to a gain of latency of up to 10x compared to baseline models.
  2. 2.
    A new inference method, Dynamic Candidate Selection (DCS), which finds an optimal number of proposals to select for the relation prediction stage, to further improve latency without compromising accuracy.
  3. 3.
    A new model, REACT, that extends previous work and our DTS architecture. REACT achieves reduced latency to just 23ms (44 FPS) while enhancing performance on standard metrics (F1@K).
Our findings demonstrate the inefficiency of current implementations of SGG models and even highlight a few drawbacks in the design of prior approaches, which can be taken into account for future work. We also want to emphasize that our method can considerably reduce the time and resources needed for training, which is a step toward more sustainable SGG implementations [33].
@@ 2 Related Work
Two-stage approaches rely on a Faster-RCNN detector with a ResNeXt-101 backbone for object detection and feature extraction [41]. Classical approaches include Neural-Motifs [44], VCTree [34], Transformer [35], GPS-Net [26] or PE-NET [46]. Faster-RCNN is not optimal for real-time constraints because of the size and complexity of the backbone. In recent work, Jin et al. [11] introduced an approach for real-time SGG based on contextual information. In contrast to other approaches, this work does not use visual features to infer relations. Instead, it uses the object’s bounding box coordinates predicted by a YOLOV5 [14] model to learn the correlation between pairs of coordinates and predicate classes. However, the reliance on spatial features only will neglect fine-grained relations such as the difference between $eating$ or $drinking$. Another work [13] proposed to achieve real-time SGG by leveraging a Relation-aware YOLO structure (RYOLO). The approach involves predicting the coordinates of relations by adding new anchors on the feature maps of YOLO. Then, matching is done with spatially proximate objects detected. However, such an approach leads to poor performance in RelPred, with only 3.1% in the mR@100 metric. This can be due to the absence of contextual information from the subject and object during relation prediction [44]. In contrast, we propose to use the features extracted by a YOLO model to generate representations for the subject and object nodes, making the first stage of the architecture more efficient without compromising performance on the second stage.
Recently, the paradigm of SGG has been tackled as a one-stage architecture, learning to represent both objects and relations at the same time. For instance, Sparse-RCNN [36] employs a strategy of triplet querying to generate object proposals and relations. Once visual features have been extracted from a CNN backbone, boxes and relations are decoded altogether using a cascade-RCNN scheme. RelationTransformer (RelTr) [3] and SGTR [21] are two approaches based on the DETR architecture [1]. Both approaches generate a sparse set of relations between $<subject,object>$ pairs, in contrast to traditional two-stage approaches which will compute probabilities for all possible relations (dense predictions). Inspired by sparse approaches in one-stage SGG, we aim to reduce the computational bottleneck of two-stage SGG by selecting only relevant object pairs for the relation prediction stage. Lately, Im et al. introduced EGTR [10], a new one-stage approach based on a Deformable-DETR architecture, more efficient than classical DETR for the OD task, paired with a graph relation matching process for RelPred. Authors of RelTr and EGTR have compared their work with two-stage approaches such as Motifs [44] and VCTree [34] in terms of latency. The comparison shows a clear disadvantage for any two-stage approaches. This may be biased by the use of Faster-RCNN for OD in two-stage approaches. We strongly believe that lower latency and better accuracy can be attained by modifying two-stage approaches to incorporate real-time object detectors such as YOLO in place of Faster-RCNN.
@@
Figure 2: Top: the PE-NET architecture [46]. Bottom: our proposed REACT model with (left) Decoupled Two-Stage (DTS) + YOLO for efficient object detection and feature extraction ; (right) our proposed Efficient Prototype Learning model. [- - -] represent modified or added components, $\bigotimes$ denotes element-wise concatenation.
@@ 3 Investigating Bottlenecks in SGG
In Figure 2 (top row) we display as current the PE-NET model [46]: a state-of-the-art two-stage SGG approach. The feature extraction (stage 1 Figure 2(a)) as well as the graph prediction (module 4, Figure 2(b)) is common to all two-stage architectures [34, 35, 6, 19, 44, 26]. In Stage 1, visual features are extracted by a ResNeXt-101 backbone (Figure 2(a) (a)), and a Feature Pyramid Network (FPN) [25] is used to refine the features at different scales (Figure 2(a) (b)). Then, a Region Proposal Network (RPN) [32] generates proposals (Figure 2(a) (c)) which are used to align three different types of features with corresponding image regions: union features, spatial features, and visual features (Figure 2(a) (d), (e), and (f)). Union and spatial features are aggregated for every possible pair of proposals and will serve as the initial edge representation in Stage 2 (right) which is the relation prediction stage (Figure 2(b)). A classifier extracts the class labels of each proposal (Figure 2(b) (h)) and retrieves corresponding linguistic features from pre-trained models (GloVe word embeddings [31]). Then, the visual features are upsampled (Figure 2(b) (i)) and combined with the textual features to form the node representation through semantic fusion (Figure 2(b) (j)). Then, prototype learning takes place (Figure 2(b) (k)). Finally, object labels and predicate labels are decoded to form the final graph representation (Figure 2(b) (l), (m), and (n)). This architecture is complex and may not be efficient for retaining the OD performance of the backbone detector and providing low latency for real-time applications. In the following, we present in more detail the bottlenecks related to the Object Detection Accuracy, Proposal Candidate Selection, and the latency of the model.
3.1 Object Detection Accuracy
@@
As we can see in Figure 2(b), the strategy employed in the two-stage SGG is to decode object labels twice during the relation prediction stage: one time for extracting text features (Figure 2(b) (h)) and one time to compute the final predictions (Figure 2(b) (m)). The predictions obtained from this last step depend on each model’s context learning strategy [34, 44, 26, 35]. This architecture design is (1) redundant as similar operations are performed twice and (2) can introduce significant differences in object detection accuracy in contrast to the original Faster-RCNN implementation.
In Table 1, we draw a comparison of the performance of the Faster-RCNN model equipped with a ResNeXt-101 feature extractor on the PSG dataset with different relation prediction heads. We can observe that the performance in OD drops significantly, from 1 point for Motifs-TDE [44, 35] to almost 13 points in mAP@50 for GPS-NET [26] after the relation prediction stage.
Table 1: Performance on Object Detection of Faster-RCNN before and after the relation prediction stage with different SGG models on the PSG dataset [42].
Backbone Relation Head mAP@50
Faster-RCNN - 36.4
Neural-Motifs [44] 35.9
PE-NET [46] 35.4
Transformer [35] 34.6
VC-Tree [34] 32.58
GPS-NET [26] 31.55
These results confirm a true dependency between both stages of the pipeline, which seems to hinder the performance of models in the OD task. Using the Faster-RCNN detector for feature extraction and object detection may not be efficient for low latency and high accuracy. In fact, real-time detectors such as YOLOV8 are today largely outperforming Faster-RCNN [15]. However, simply replacing Faster-RCNN with YOLO in the architecture of Two-Stage SGG is not possible because YOLO does not possess an RPN or ROI Align component.
3.2 Proposal Candidate Selection
Subsequently, performing Non-Maximum Suppression (NMS) (Figure 2(b) (n)) after the context learning is inefficient because a lot of false positive or irrelevant proposals will be selected as node candidates during features refinement and context learning. To alleviate the lack of NMS before stage 1, a fixed set of the top proposals (i.e. 80 or 100) are traditionally considered valid node candidates for stage 2. The idea here is that by sampling a large number of proposals per image, it is easier to find valid pairs and relations. However, this considerably increases the computational complexity required to generate the graph, which is a bottleneck for real-time applications. Having a large number of proposals as candidates can make the final graph less informative, as some important objects can be left out if they do not possess at least one high confidence relation. To alleviate this issue, prior works [44, 35, 19] compute the final ranking of relations with the following score formula:
$\theta_{rel}=\theta_{obj}*\theta_{pred}*\theta_{subj}$ (1)
[…]
3.3 Latency of the PE-NET model
@@
Lastly, we investigate the bottlenecks in latency for the relation prediction stage (see Figure 2(b)) of the PE-NET model [46]. In PE-NET, node prototypes are formed using a combination of linguistic features and visual features for each independent proposal (see Figure 2(b) (j)). Edge features are computed using the concatenation of the spatial and union features (see Figure 2(b) (g)). During prototype learning, prototype representations of the subject and object are combined and added to the edge features to form the relation prototype. Finally, a gate mechanism is employed to select relevant features from the multi-modal fusion. The use of both union and spatial features for the edge representation represents a consequent amount of overload for the model. In Figure 2(a) we see that another step of ROI Align is performed on the union of subject and object boxes in step (d). However, a lot of visual information is already present in the object and subject features (see Figure 2(a) (f)) which will be fused with edge features during prototype learning. In the original work [46], this design is not justified. In addition, during edge and node feature refinement, all features are traditionally upsampled to a fixed size of 4096 (see Figure 2(b) (g) and Figure 2(b) (i)). However, larger feature sizes can significantly augment the latency of the model. Therefore, we aim to investigate the usage and size of each feature type to find an optimal trade-off between latency and performance.

### v3

@@ Abstract
Scene Graph Generation (SGG) is a task that encodes visual relationships between objects in images as graph structures. SGG shows significant promise as a foundational component for downstream tasks, such as reasoning for embodied agents. To enable real-time applications, SGG must address the trade-off between performance and inference speed. However, current methods tend to focus on one of the following: (1) improving relation prediction accuracy, (2) enhancing object detection accuracy, or (3) reducing latency, without aiming to balance all three objectives simultaneously. To address this limitation, we propose the Real-time Efficiency and Accuracy Compromise for Tradeoffs in Scene Graph Generation (REACT) architecture, which achieves the highest inference speed among existing SGG models, improving object detection accuracy without sacrificing relation prediction performance. Compared to state-of-the-art approaches, REACT is 2.7 times faster and improves object detection accuracy by 58%. Furthermore, our proposal significantly reduces model size, with an average of 5.5x fewer parameters. The code is available at https://github.com/Maelic/SGG-Benchmark.
$\dagger$ Corresponding authors.
Figure 1: Performance comparison between our approach and previous work, in terms of performance (F1@K metric) and Latency (ms) on the PSG dataset [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu]. Points are scaled by the number of parameters (smaller is better).
@@ 1 Introduction
Scene Graph Generation (SGG) is the task of generating a structured graph representation from visual inputs. Given an image, visual regions corresponding to objects of interest are extracted, and relations between these objects are predicted in the form of <subject, predicate, object> triplets. The set of these triplets forms a directed acyclic graph which can be interpreted as a structured representation of the scene. This task has garnered increasing attention recently thanks to its utility in various downstream applications, such as Visual Question Answering (VQA) [Hudson and Manning(2019), Damodaran et al.(2021)Damodaran, Chakravarthy, Kumar, Umapathy, Mitamura, Nakashima, Garcia, and Chu, Liang et al.(2021)Liang, Jiang, and Liu], Image Captioning [Yang et al.(2022b)Yang, Liu, and Wang, Wang et al.(2019)Wang, Beck, and Cohn, Nguyen et al.(2021)Nguyen, Tripathi, Du, Guha, and Nguyen], or Reasoning of an embodied agent [Gadre et al.(2022)Gadre, Ehsani, Song, and Mottaghi, Li et al.(2022b)Li, Guo, Liu, and Sun, Neau et al.(2023a)Neau, Santos, Bosser, and Buche]. However, we observe a significant imbalance between the progress in the SGG task itself and its adoption in downstream tasks [Chang et al.(2023)Chang, Ren, Xu, Li, Chen, and Hauptmann]. One of the reasons is the lack of low-cost approaches that could power applications with real-time constraints. Because SGG is a task that combines Object Detection (OD) and Relation Prediction (RelPred), a good trade-off between accuracy in these two parameters is also required. In this work, we explore the bottlenecks in SGG to find an optimal combination of OD accuracy, RelPred accuracy, and latency. SGG methods can be either Two-Stage (TS) or One-Stage (OS) approaches. The former uses a combination of an object detector and a relation predictor to model objects and relations in a sequence. The latter uses a single-stage pipeline to infer relations and object proposals directly from the image features. As a result, one-stage approaches are often more efficient for real-time processing [Im et al.(2024b)Im, Nam, Park, Lee, and Park]. However, they fall short in accuracy in the OD task [Li et al.(2022a)Li, Zhang, and He, Im et al.(2024a)Im, Nam, Park, Lee, and Park].   In this work, we investigate the bottlenecks of the two-stage architecture in terms of OD accuracy and latency. The structural dependencies between blocks of the architecture are reviewed as a Decoupled Two-Stage SGG (DTS-SGG), which allows maintaining the performance of the pre-trained detector, providing better control of the information flow (i.e., detected object proposals) between the two stages of the architecture. This approach reduces the overall latency, as the complexity of relation prediction increases exponentially with the number of input objects. To select an optimal amount of proposals at inference time, we define a new algorithm, the Dynamic Candidate Selection (DCS). DCS uses evaluation results to find the best trade-off between latency and accuracy at test time. Thanks to the DTS architecture, we can switch the dependency on a Faster-RCNN model for any real-time object detectors, such as one of the YOLO family (YOLOV8/9/10 [Jocher et al.(2023)Jocher, Chaurasia, and Qiu, Wang et al.(2024b)Wang, Yeh, and Liao, Wang et al.(2024a)Wang, Chen, Liu, Chen, Lin, Han, and Ding]). Recent state-of-the-art SGG models such as PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song] use a combination of different feature sources, such as union pair features [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] or textual features of the subject and object classes [Lu et al.(2016)Lu, Krishna, Bernstein, and Fei-Fei]. This results in a significant amount of processing to gather and refine each type of feature altogether. By removing the dependency on sub-optimal features, we can reduce the parameter count of the model and lower the latency. By coupling this strategy to the DTS architecture, we propose a new model for SGG, the Real-Time Efficiency and Accuracy Compromise for Tradeoffs in SGG (REACT). Thus, the main contributions of this work are: (i) a new architecture for SGG, the Decoupled Two-Stage (DTS); (ii) a new inference method, Dynamic Candidate Selection (DCS), which finds an optimal number of proposals to select for the relation prediction stage, to further improve latency without compromising accuracy; (iii) a new model, REACT, that extends previous work and our DTS architecture, reducing the latency while enhancing performance on standard metrics, see Figure˜1.
@@ 2 Related Work
Two-stage SGG approaches typically use Faster-RCNN with a VGG-16 [Xu et al.(2017)Xu, Zhu, Choy, and Fei-Fei, Yang et al.(2018)Yang, Lu, Lee, Batra, and Parikh] or ResNeXt-101 backbone for object detection and feature extraction, as seen in Neural-Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi], VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], Transformer [Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang], GPS-Net [Lin et al.(2020)Lin, Ding, Zeng, and Tao], and PE-NET [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song]. In these approaches, object classes probabilities given by Faster-RCNN are refined during relationship learning, via a dedicated classifier [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao] or by aggregating features from neighboring regions [Yang et al.(2018)Yang, Lu, Lee, Batra, and Parikh]. However, this design can hinder the performance of SGG models on the task of Object Detection. In our work, we choose to effectively Decoupled the Object Detector and Relationship Predictor by removing the dedicated classifier and keeping original class probabilities. In addition, Faster-RCNN’s large backbone limits real-time use. Recent work [Jin et al.(2023a)Jin, Guo, Meng, Zhu, Xi, Wang, Mu, and Song] proposed a real-time SGG method using contextual cues, relying on YOLOv5 [Jocher(2020)] bounding boxes to learn correlations between box coordinates and predicates—excluding visual features. However, the method’s reliance on spatial features will neglect fine-grained relations such as the difference between eatingeating or drinkingdrinking. Another approach [Jin et al.(2023c)Jin, Wang, Zhu, Xi, Meng, Mu, and Song] uses a Relation-aware YOLO (RYOLO) to predict relation coordinates via new anchors, matched to nearby objects. Yet, lacking subject-object context leads to poor RelPred performance (3.1% mR@100) [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi]. In contrast, we propose to use the features extracted by a YOLO model to generate representations for the subject and object nodes, making the first stage of the architecture more efficient without compromising performance on the second stage.
One-stage SGG approaches jointly learn object and relation representations. Sparse-RCNN [Teng and Wang(2022a)] uses triplet queries to produce objects and relations, decoded via cascade-RCNN after CNN feature extraction. RelTr [Cong et al.(2022a)Cong, Yang, and Rosenhahn] and SGTR [Li et al.(2024)Li, Zhang, and He], built on DETR [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], generate sparse relation sets, unlike dense two-stage predictions. More recently, EGTR [Im et al.(2024b)Im, Nam, Park, Lee, and Park] builds upon Deformable-DETR for efficient object detection, paired with graph-based relation matching. RelTr and EGTR report lower latency than two-stage models like Motifs [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi] and VCTree [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu], though these comparisons often rely on Faster-RCNN. As highlighted in recent work [Li et al.(2024)Li, Zhang, and He], learning both the object and relation representations at the same time can make the performance of one-stage models in OD sub-optimal in comparison to two-stage approaches. We argue that replacing Faster-RCNN with YOLO in two-stage pipelines can achieve both lower latency than two-stage approaches and higher OD accuracy than one-stage approaches, as we shall see in this paper.
@@
Figure 2: Top: the PE-NET architecture [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song]. Bottom: our proposed REACT model with (left) Decoupled Two-Stage (DTS) + YOLO for efficient object detection and feature extraction ; (right) our proposed Efficient Prototype Learning model. [- - -] represent modified or added components, $\bigotimes$ denotes element-wise concatenation.
@@ 3 Investigating Bottlenecks in SGG
Figure˜2 (top row) is a representation of the PE-NET model [Zheng et al.(2023)Zheng, Lyu, Gao, Dai, and Song]: a state-of-the-art two-stage SGG approach. The feature extraction (stage 1 Figure˜2(a)) as well as the graph prediction (module 4, Figure˜2(b)) are common to all two-stage architectures [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Dong et al.(2022)Dong, Gan, Song, Wu, Cheng, and Nie, Li et al.(2021)Li, Zhang, Wan, and He, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao]. In Stage 1, visual features are extracted by a ResNeXt-101 backbone (Figure˜2(a) (a)), and a Feature Pyramid Network (FPN) [Lin et al.(2017)Lin, Dollár, Girshick, He, Hariharan, and Belongie] is used to refine the features at different scales (Figure˜2(a) (b)). Then, a Region Proposal Network (RPN) [Ren et al.(2015)Ren, He, Girshick, and Sun] generates proposals (Figure˜2(a) (c)) which are used to align three different types of features with corresponding image regions: union features, spatial features, and visual features (Figure˜2(a) (d), (e), and (f)) using ROI Align [Ren et al.(2015)Ren, He, Girshick, and Sun]. Union and spatial features are aggregated for every possible pair of proposals and will serve as the initial edge representation to the relation prediction stage (Stage 2, Figure˜2(b)). A classifier extracts the class labels of each proposal (Figure˜2(b) (h)) and retrieves corresponding linguistic features from pre-trained text embeddings (GloVe [Pennington et al.(2014)Pennington, Socher, and Manning]). Then, the visual features are combined with the textual features to form the node representation through semantic fusion (Figure˜2(b) (i), (j)). Then, prototype learning takes place (Figure˜2(b) (k)). Finally, object and predicate labels are decoded to form the final graph representation (Figure˜2(b) (l), (m), and (n)). Below, we discuss the main bottlenecks of this architecture for Latency and Object Detection accuracy.
@@
Object Detection Accuracy. The strategy employed in the two-stage SGG is to decode object labels twice during the relation prediction stage: one time for extracting text features (Figure˜2(b) (h)) and one time to compute the final predictions (Figure˜2(b) (m)). The predictions obtained from this last step depend on each model’s context learning strategy [Tang et al.(2019)Tang, Zhang, Wu, Luo, and Liu, Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Lin et al.(2020)Lin, Ding, Zeng, and Tao, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang]. This architecture design is (i) redundant as similar operations are performed twice, and (ii) can introduce significant differences in object detection accuracy in contrast to the original Faster-RCNN implementation. Table˜1 draws a comparison of the performance of the Faster-RCNN model equipped with a ResNeXt-101 feature extractor on the PSG dataset with different relation prediction heads. We can observe that the performance in OD drops significantly, from 1 point wrt Motifs-TDE [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang] to almost 13 points in mAP@50 wrt GPS-NET [Lin et al.(2020)Lin, Ding, Zeng, and Tao] after the relation prediction stage. These results demonstrate a clear dependency between the two stages of the pipeline, which negatively impacts model performance on the object detection (OD) task. A second consideration given these results is the low performance of Faster-RCNN. Recent state-of-the-art detectors such as YOLOV8/9/10 may be more appropriate for the task. In contrast to Faster-RCNN, the YOLO architecture is not designed for Feature Extraction. This is the main reason why the Faster-RCNN architecture is still widely used for SGG.
Table 1: OD accuracy for two-stage models compared to Faster-RCNN, PSG dataset [Yang et al.(2022a)Yang, Ang, Guo, Zhou, Zhang, and Liu].
Proposal Candidate Selection. In the TS architecture, Non-Maximum Suppression (NMS) for object proposals (Figure˜2(b) (n)) is performed in stage 2. To alleviate the lack of NMS in stage 1, a fixed set of the top proposals is traditionally considered valid node candidates for stage 2. The idea here is that by sampling a large number of proposals per image, it is easier to find valid pairs and relations. However, this increases the computational complexity of the graph learning. Also, the number of top proposals to select (traditionally 80 or 100) is set arbitrarily. Recent work [Zellers et al.(2018)Zellers, Yatskar, Thomson, and Choi, Tang et al.(2020)Tang, Niu, Huang, Shi, and Zhang, Li et al.(2021)Li, Zhang, Wan, and He] computes the final ranking of relations with the following score formula: $\theta_{rel}=\theta_{obj}*\theta_{pred}*\theta_{subj},$ where $\theta_{pred}$ is the confidence score of a predicate, given <subject,object><subject,object> pair as candidate; and, $\theta_{obj},\theta_{subj}$ are the respective confidence score of the object detector. This formulation assigns greater weight to the confidence scores of the subject and object compared to the predicate. As a result, many low-confidence proposals can be filtered out prior to the relation prediction stage with minimal impact on the model’s overall accuracy.
@@
Latency of the PE-NET model. In PE-NET, node prototypes are formed using a combination of linguistic features and visual features (see Figure˜2(b) (j)). Edge features are computed using the concatenation of the spatial and union features (see Figure˜2(b) (g)). During prototype learning, prototype representations of the subject and object are combined and added to the edge features to form the relation prototype. Finally, a gate mechanism is employed to select relevant features from the multi-modal fusion. The use of both union and spatial features represents a considerable amount of overload for the model. In fact, another step of ROI Align is performed on the union of subject and object boxes (Figure˜2(a) step (d)). However, visual information is already present in the object and subject features (see Figure˜2(a) (f)), which will be fused with edge features during prototype learning. In addition, during edge and node feature refinement, all features are traditionally upsampled to a fixed size of 4096 (see Figure˜2(b) (g) and Figure˜2(b) (i)). However, larger feature sizes can significantly augment the latency of the model. Therefore, we aim to investigate the usage and size of each feature type to find an optimal trade-off between latency and performance.